PHYS 1111L Part B Analyzing Velocity-Time Graphs Velocity (em/sec) 15 Time (sec) 10 -5 -15 Data Table for Part B- Analyzing Velocity-Time Graphs Ax X initial X final Interval | v 0-2 sec 2-4 sec 4-6 sec 6-8 sec 8-10 sec a average V final
Answer #1

The graph given above is a velocity vs time graph from 0 < t < 10s

for 0 - 2s, vinitial = vfinal = vaverage = 0 m/s

aaverage = 0 m/s2

and so displacement Delta x and final position is also 0 m.

for 2-4s, vinitial = 0 m/s

vfinal = 15 m/s

therefore, vaverage = (0+15/2) = 7.5 m/s

=> aaverage = 7.5/(4 - 2) = 3.75 m/s2

use vfinal = vinitial + at

=> 15 = 0 + a(4 - 2)

=> a = 7.5 m/s2

therefore, ur = linitialt +-at2 = 15m ,T U

xinitial = 0 m and so xfinal = xinitial + Delta x = 0 + 15 = 15 m.

Repeat for other time segments.